# LMI Optimization Based Multirate Steady-State Kalman Filter Design

### v1

@@ Abstract
This paper presents an LMI-based design framework for multirate steady-state Kalman filters in systems with sensors operating at different sampling rates. The multirate system is formulated as a periodic time-varying system, where the Kalman gains converge to periodic steady-state values that repeat every frame period. Cyclic reformulation transforms this into a time-invariant problem; however, the resulting measurement noise covariance becomes semidefinite rather than positive definite, preventing direct application of standard Riccati equation methods. We address this through a dual LQR formulation with LMI optimization that naturally handles semidefinite covariances. The framework enables multi-objective design, supporting pole placement for guaranteed convergence rates and mixed $H_{2}$$l_{2}$-induced norm design for balancing average and worst-case performance. Numerical validation using an automotive navigation system with GPS and wheel speed sensors demonstrates that the proposed filter achieves estimation errors well below raw measurement noise levels.
@@ 1 Introduction
Several approaches address multirate Kalman filtering. The lifting technique [5, 6, 23, 16, 7] transforms multirate systems into equivalent time-invariant systems at the frame period by augmenting input and output vectors. While preserving state dimension, lifting requires matrix products such as $A^{N}$ and operates in open-loop between frame periods, limiting inter-sample accuracy. In contrast, the cyclic reformulation [19] employed in this paper maintains the original state dimension within each block and explicitly represents periodicity through cyclic shift of state variables, enabling LMI-based optimization that naturally handles semidefinite measurement noise covariances. For systems with measurement delays, pseudo-measurement approaches [8, 9] reconstruct current-time observations from delayed measurements and predicted states. Asynchronous filtering [10, 11] processes measurements upon arrival but needs careful update handling.
The theoretical foundation of periodic systems and periodic Riccati equations has been extensively studied. Bittanti, Colaneri, and their collaborators established fundamental results on discrete-time periodic Riccati equations (DPRE), including existence, uniqueness, and convergence properties of periodic solutions [24, 25]. Varga [26] developed numerically reliable algorithms for solving both continuous-time and discrete-time periodic Riccati equations using periodic Schur decomposition. De Souza [27] addressed the periodic Riccati difference equation specifically for optimal filtering of linear periodic discrete-time systems.
@@
The main contributions of this paper are as follows. First, we identify that the cyclic measurement noise covariance $\check{R}$ is positive semidefinite but not positive definite in realistic multirate systems, and present an LMI-based approach via dual LQR formulation that handles this naturally. Second, we develop LMI-based extensions for multi-objective design, including pole placement constraints for guaranteed convergence rates and Kalman filter design with $l_{2}$-induced norm constraints [22, 21] for balancing average and worst-case performance. Third, we demonstrate the effectiveness of the proposed methods through automotive state estimation simulations with GPS and wheel speed sensors operating at different rates. The LMI framework also provides additional benefits: stability guarantees through the established connection between cyclic system stability and the monodromy matrix, and straightforward extension to weighted optimal estimation by modifying the objective function.
@@ 2.1 Standard Kalman Filter
In this paper, we adopt the predictor form [14], which is standard in observer-based control design. The corresponding gain $L_{ss}=AK_{ss}$ yields the estimator:
@@ 3.1 Problem Formulation
Suppose the system is equipped with multiple sensors operating at different sampling rates. Let $C\in\mathbb{R}^{q\times n}$ denote the full measurement matrix, where $q$ is the total number of scalar measurements. Since not all sensors provide measurements at every sampling instant, we introduce a periodic selection matrix $S_{k}\in\mathbb{R}^{q\times q}$ [21, 20] to represent measurement availability. The measurement equation for the multirate system is:
@@ 3.2 Cyclic Reformulation
To transform the periodic system (16)–(17) into a time-invariant form, we apply cyclic reformulation [19], following its extension to multirate systems in [22, 20]. For a signal $\xi(k)$, define its cycled representation $\check{\xi}(k)$ by placing the signal in the $(k\bmod N+1)$-th block position with zeros elsewhere. Specifically, for $k\bmod N=j$ $(j=0,1,\ldots,N-1)$:
@@ 3.3 Observability Analysis
The observability of the cyclic system is crucial for the existence of a stable Kalman filter. The observability matrix for the cyclic system is:
@@
If the pair $(A,S_{k}C)$ is observable over one period (i.e., the periodic observability matrix has full rank), then the cyclic system $(\check{A},\check{C})$ is observable.
@@ 3.4 Challenges with Standard DARE Approach
$\check{L}=\check{P}\check{C}^{T}(\check{C}\check{P}\check{C}^{T}+\check{R})^{-1}$ (34)
@@ 3.5 Stability Analysis
Corollary 3.1
When the LMI optimization (42) in Section 4 yields a feasible solution with $\check{X}\succ 0$, the resulting filter is guaranteed to be stable, i.e., $\rho(\check{A}-\check{L}\check{C})<1$.
This stability guarantee is inherent in the LMI formulation, as the constraint (39) implicitly enforces Schur stability through the Lyapunov inequality structure. The complete design procedure is presented in Section 4.
@@ 4.1 LMI-based Filter Design
4.1.1 Dual LQR Formulation
The key insight is to reformulate the Kalman filtering problem as a dual Linear Quadratic Regulator (LQR) problem [14]. The Kalman filter for the system $(\check{A},\check{C},\check{Q},\check{R})$ is equivalent to the LQR problem for the dual system:
$\check{\zeta}(k+1)$ $=\check{A}^{T}\check{\zeta}(k)+\check{C}^{T}\check{u}(k)$ (38)
with state weighting $\check{Q}$ and input weighting $\check{R}$. The optimal LQR gain directly yields the Kalman gain through this duality, enabling the application of LMI-based methods [15] to the filtering problem while naturally accommodating the semidefinite $\check{R}$.
@@ 4.1.2 LMI Formulation
Define the decision variables:
  • $\check{X}\in\mathbb{R}^{Nn\times Nn}$: $\check{X}\succ 0$ (Lyapunov matrix)
  • $\check{Y}\in\mathbb{R}^{Nn\times Nq}$: $\check{Y}=-\check{X}\check{L}$ where $\check{L}$ is the Kalman gain
The LMI formulation consists of the following constraints:
LMI #1: Stability and Performance (DARE-based LMI)
@@
where $\check{Q}^{1/2}$ has the same cyclic structure as $\check{A}$ with $Q^{1/2}$ placed in the cyclic positions, and $\check{R}^{1/2}$ is the block-diagonal matrix with $S_{k}R^{1/2}$ on the diagonal blocks.
@@ 4.1.4 Kalman Gain Extraction
Note: To validate the proposed LMI solution $\check{X}^{-1}\check{Y}$, we conducted numerical experiments with $\check{R}\succ 0$, where the standard DARE solution exists. Let $\check{K}_{\text{Ric}}$ denote the filter gain obtained from the DARE. The predictor-form gain is then $\check{L}_{\text{Ric}}=\check{A}\check{K}_{\text{Ric}}$. Numerical experiments confirm $\|-\check{X}^{-1}\check{Y}-\check{L}_{\text{Ric}}\|_{F}<10^{-6}$.
@@ 5.0.1 Pole Placement Constraints
To guarantee a minimum convergence rate for the estimation error, we can constrain the eigenvalues of the error dynamics matrix $\check{A}-\check{L}\check{C}$ to lie within a specified region. For discrete-time systems, a common choice is the disk region:
@@ 5.0.2 Kalman Filter with $l_{2}$-induced Norm Constraint
The LMI (53) admits a scaling degree of freedom: replacing $\check{C}_{z}$ with $\alpha\check{C}_{z}$ and $\gamma^{2}$ with $\gamma^{2}/\alpha^{2}$ for any $\alpha>0$ yields an equivalent $l_{2}$-induced norm condition. This scaling is useful when the (2,2) block $\check{X}-\check{C}_{z}^{T}\check{C}_{z}$ causes infeasibility; choosing $\alpha<1$ relaxes the constraint $\check{X}\succ\check{C}_{z}^{T}\check{C}_{z}$, enabling a common Lyapunov matrix $\check{X}$ to satisfy both the Kalman filter LMI (39) and the $l_{2}$-induced norm constraint simultaneously.
@@ 5.0.3 Combined Multi-objective Design
where $\check{\Gamma}\succ 0$ is a state weighting matrix.
@@ 6 Numerical Example: Multirate Automotive State Estimation
To demonstrate the effectiveness of the proposed multirate Kalman filter design method, we consider an automotive navigation system as a practical application. This example illustrates how the cyclic reformulation framework can be applied to a real-world multirate sensing scenario where GPS and wheel speed sensors operate at different sampling rates.
@@ 6.1 System Configuration
We consider an automotive navigation system with the following specifications:
@@ 6.4 Simulation Results
A simulation of 200 time steps (20 seconds) was conducted with:
  • Initial state: $x(0)=[0,5,0]^{T}$ (position 0 m, velocity 5 m/s)
  • Control input: $u(k)=0.5\sin(0.05k)$ (sinusoidal acceleration)
  • Process and measurement noise: zero-mean Gaussian with covariances $Q$ and $R$
The estimation performance is evaluated using root mean square error (RMSE):
$\text{RMSE}_{\text{pos}}$ $=0.600\text{ m}$ (67)
$\text{RMSE}_{\text{vel}}$ $=0.268\text{ m/s}$ (68)
$\text{RMSE}_{\text{acc}}$ $=1.165\text{ m/s}^{2}$ (69)
The multirate Kalman filter successfully tracks the vehicle motion, with estimation errors significantly reduced compared to the raw sensor measurements. Figs. 13 show the estimation results for position, velocity, and acceleration, respectively. Position estimation errors are largest just before GPS updates (Fig. 1) and are corrected sharply when GPS measurements become available.

### v3

@@ Abstract
This paper presents an LMI-based design framework for multirate steady-state Kalman filters in systems with sensors operating at different sampling rates. The multirate system is formulated as a periodic time-varying system, where the Kalman gains converge to periodic steady-state values that repeat every frame period. Cyclic reformulation transforms this into a time-invariant problem; however, the resulting measurement noise covariance becomes semidefinite rather than positive definite, preventing direct application of standard Riccati equation methods. I address this through a dual LQR formulation with LMI optimization that naturally handles semidefinite covariances. The framework enables multi-objective design, supporting pole placement for guaranteed convergence rates and $l_{2}$-induced norm constraints for balancing average and worst-case performance. Numerical validation using an automotive navigation system with GPS and wheel speed sensors, including Monte Carlo simulation with 500 independent noise realizations, demonstrates that the proposed filter achieves a position RMSE well below the GPS noise level through effective multirate sensor fusion, and that the LMI solution provides valid upper bounds on the estimation error covariance.
@@ 1 Introduction
Several approaches address multirate Kalman filtering. The lifting technique [5, 22, 15, 6] transforms multirate systems into equivalent time-invariant systems at the frame period by augmenting input and output vectors. While preserving state dimension, lifting requires matrix products such as $A^{N}$ and operates in open-loop between frame periods, limiting inter-sample accuracy. In contrast, the cyclic reformulation [18] employed in this paper maintains the original state dimension within each block and explicitly represents periodicity through cyclic shift of state variables, enabling LMI-based optimization that naturally handles semidefinite measurement noise covariances. For systems with measurement delays, pseudo-measurement approaches [7, 8] reconstruct current-time observations from delayed measurements and predicted states. Asynchronous filtering [9, 10] processes measurements upon arrival but needs careful update handling.
The theoretical foundation of periodic systems and periodic Riccati equations has been extensively studied. Bittanti, Colaneri, and their collaborators established fundamental results on discrete-time periodic Riccati equations (DPRE), including existence, uniqueness, and convergence properties of periodic solutions [23, 24]. Varga [25] developed numerically reliable algorithms for solving both continuous-time and discrete-time periodic Riccati equations using periodic Schur decomposition. De Souza [26] addressed the periodic Riccati difference equation specifically for optimal filtering of linear periodic discrete-time systems. For systems with singular measurement noise covariance, Nikoukhah et al. [30] studied Kalman filtering and generalized Riccati equations in the descriptor system framework.
@@
The main contributions of this paper are as follows. First, I identify that the cyclic measurement noise covariance $\check{R}$ is positive semidefinite but not positive definite in realistic multirate systems, and present an LMI-based approach via dual LQR formulation that handles this naturally. Second, I develop LMI-based extensions for multi-objective design, including pole placement constraints for guaranteed convergence rates and Kalman filter design with $l_{2}$-induced norm constraints [21, 20] for balancing average and worst-case performance. Third, I demonstrate the effectiveness of the proposed methods through automotive state estimation simulations with GPS and wheel speed sensors operating at different rates. The LMI framework also provides additional benefits: stability guarantees through the established connection between cyclic system stability and the monodromy matrix, and straightforward extension to weighted optimal estimation by modifying the objective function.
@@ 2.1 Standard Kalman Filter
In this paper, I adopt the predictor form [13], which is standard in observer-based control design. The filter-form gain $K_{ss}$ is converted to the predictor-form gain $L_{ss}=AK_{ss}$ by premultiplying by $A$, so that the measurement correction is incorporated into the one-step-ahead state prediction. The corresponding estimator is:
@@ 3.1 Problem Formulation
Suppose the system is equipped with multiple sensors operating at different sampling rates. Let $C\in\mathbb{R}^{q\times n}$ denote the full measurement matrix, where $q$ is the total number of scalar measurements. Since not all sensors provide measurements at every sampling instant, I introduce a periodic selection matrix $S_{k}\in\mathbb{R}^{q\times q}$ [20, 19] to represent measurement availability. The measurement equation for the multirate system is:
@@ 3.2 Cyclic Reformulation
To transform the periodic system (16)–(17) into a time-invariant form, I apply cyclic reformulation [18], following its extension to multirate systems in [21, 19]. For a signal $\xi(k)$, define its cycled representation $\check{\xi}(k)$ by placing the signal in the $(k\bmod N+1)$-th block position with zeros elsewhere. Specifically, for $k\bmod N=j$ $(j=0,1,\ldots,N-1)$:
@@ 3.3 Observability Analysis
The observability of the cyclic system is crucial for the existence of a stable Kalman filter. The observability matrix for the $Nn$-dimensional cyclic system $(\check{A},\check{C})$ is:
@@
The cyclic system $(\check{A},\check{C})$ is observable if and only if $\mathrm{rank}(\mathcal{O})=Nn$. Since the system matrices $A$, $S_{k}$, and $C$ are assumed known, this condition can be verified directly by computing the rank of $\mathcal{O}$.
This is a collective observability condition over one complete period of $N$ time steps: the combination of all measurements collected over one period must provide sufficient information to reconstruct the full state. Individual time steps need not be independently observable.
Remark 3.2 (Sufficient Conditions and Detectability)
A sufficient condition for the observability of $(\check{A},\check{C})$ is that there exists some $j\in\{0,\ldots,N-1\}$ such that the pair $(S_{j}C,A^{N})$ is observable and $A$ is nonsingular [29]. In this case, the nonsingularity of $A$ propagates the observability at a single measurement time instant to all block positions in the cyclic observability matrix, establishing $\mathrm{rank}(\mathcal{O})=Nn$. For the stability of the Kalman filter, the weaker condition of detectability of $(\check{A},\check{C})$ is sufficient: all unobservable modes of the cyclic system must be Schur stable. In particular, when $A$ is Schur stable, detectability is automatically satisfied regardless of the measurement pattern.
@@ 3.4 Challenges with Standard DARE Approach
$\check{L}=\check{A}\check{P}\check{C}^{T}(\check{C}\check{P}\check{C}^{T}+\check{R})^{-1}$ (34)
@@ 3.5 Stability Analysis
The complete design procedure, including the stability guarantee of the resulting filter, is presented in Section 4.
@@ 4.1 LMI-based Filter Design
4.1.1 Error Dynamics
The estimation error $\check{e}(k)=\check{x}(k)-\check{\hat{x}}(k)$ satisfies:
For any stabilizing gain $\check{L}$, the steady-state error covariance $\check{P}_{e}=\lim_{k\to\infty}\mathbb{E}[\check{e}(k)\check{e}(k)^{T}]$ satisfies the discrete Lyapunov equation (assuming $\mathbb{E}[\check{d}_{w}\check{d}_{v}^{T}]=0$):
$\check{P}_{e}=(\check{A}-\check{L}\check{C})\check{P}_{e}(\check{A}-\check{L}\check{C})^{T}+\check{Q}+\check{L}\check{R}\check{L}^{T}$ (39)
This equation holds for an arbitrary stabilizing $\check{L}$; when the optimal gain (34) is substituted, $\check{P}_{e}$ coincides with the DARE solution $\check{P}$ in (33). Note that, unlike the DARE (33), equation (39) does not involve the inverse of $\check{C}\check{P}\check{C}^{T}+\check{R}$ and is therefore well-defined even when $\check{R}$ is only positive semidefinite.
By the duality between Kalman filtering and LQR [13], the optimal gain $\check{L}$ for (39) equals $\check{F}^{T}$, where $\check{F}$ is the LQR gain for the dual system $\check{\zeta}(k+1)=\check{A}^{T}\check{\zeta}(k)+\check{C}^{T}\check{u}(k)$ with state weight $\check{Q}$ and input weight $\check{R}$. This duality enables the application of standard LQR-LMI methods [14], as developed below.
@@ 4.1.2 LMI Formulation
Introduce the upper bound variable $\check{X}\succ 0$ satisfying $\check{P}_{e}\preceq\check{X}^{-1}$. Replacing $\check{P}_{e}$ with $\check{X}^{-1}$ on both sides of (39) gives the sufficient condition for maintaining this bound:
$\check{X}^{-1}\succeq(\check{A}-\check{L}\check{C})\check{X}^{-1}(\check{A}-\check{L}\check{C})^{T}\\ +\check{Q}^{1/2}(\check{Q}^{1/2})^{T}+\check{L}\check{R}^{1/2}(\check{R}^{1/2})^{T}\check{L}^{T}$ (40)
Define the change of variables $\check{Y}=-\check{X}\check{L}$. Pre- and post-multiplying (40) by $\check{X}$ and using $\check{X}(\check{A}-\check{L}\check{C})=\check{X}\check{A}+\check{Y}\check{C}$:
$\check{X}\succeq(\check{X}\check{A}+\check{Y}\check{C})\check{X}^{-1}(\check{X}\check{A}+\check{Y}\check{C})^{T}\\ +\check{X}\check{Q}^{1/2}(\check{Q}^{1/2})^{T}\check{X}+\check{Y}\check{R}^{1/2}(\check{R}^{1/2})^{T}\check{Y}^{T}$ (41)
The right-hand side of (41) consists of three terms, each of the form $\Phi_{i}M_{i}^{-1}\Phi_{i}^{T}$ with $M_{1}=\check{X}$, $M_{2}=I_{Nn}$, $M_{3}=I_{Nq}$. Applying the Schur complement to each term simultaneously, the matrix inequality (41) is equivalent to the following block matrix inequality.
LMI #1: Stability and Performance
@@
[…]
Each block of (42) corresponds to a term in the error dynamics (38): the (1,2) block $\check{X}\check{A}+\check{Y}\check{C}$ encodes the closed-loop dynamics $\check{A}-\check{L}\check{C}$; the (1,3) block $\check{X}\check{Q}^{1/2}$ corresponds to the process noise; and the (1,4) block $\check{Y}\check{R}^{1/2}$ corresponds to the measurement noise. The diagonal blocks $I_{Nn}$ and $I_{Nq}$ at positions (3,3) and (4,4) arise from the unit-covariance normalization of $\check{d}_{w}$ and $\check{d}_{v}$.
@@ 4.1.4 Kalman Gain Extraction
4.1.5 Stability Guarantee
Theorem 4.1 (Stability Guarantee via LMI Feasibility)
If the LMI optimization (45) yields a feasible solution with $\check{X}\succ 0$, then the resulting filter gain $\check{L}=-\check{X}^{-1}\check{Y}$ guarantees Schur stability of $\check{A}-\check{L}\check{C}$, regardless of whether $\check{R}$ is positive definite or only positive semidefinite.
Proof 2
From the LMI (42) with $\check{X}\succ 0$, applying the Schur complement with respect to the (1,1) block $\check{X}$ yields a $3\times 3$ block inequality. Its (1,1) block gives:
$(\check{X}\check{A}+\check{Y}\check{C})\check{X}^{-1}(\check{X}\check{A}+\check{Y}\check{C})^{T}\preceq\check{X}$ (49)
where the noise-related terms have been dropped since they are positive semidefinite. Pre- and post-multiplying (49) by $\check{X}^{-1}$ and substituting $\check{Y}=-\check{X}\check{L}$:
$(\check{A}-\check{L}\check{C})\check{X}^{-1}(\check{A}-\check{L}\check{C})^{T}\preceq\check{X}^{-1}$ (50)
This is a discrete Lyapunov inequality with $\check{X}^{-1}\succ 0$, which is a necessary and sufficient condition for $\check{A}-\check{L}\check{C}$ to be Schur stable. The condition depends solely on the closed-loop matrix $\check{A}-\check{L}\check{C}$ and the Lyapunov matrix $\check{X}^{-1}$, and does not involve $\check{R}$.
Remark 4.1 (Roles of Observability and Semidefinite $\check{R}$)
The rank-deficiency of $\check{R}$ affects the filter design computation: the standard DARE (33) requires inversion of $\check{C}\check{P}\check{C}^{T}+\check{R}$, which fails when $\check{R}$ is singular. The LMI formulation circumvents this entirely. Once a feasible LMI solution is obtained, the closed-loop stability follows from Theorem 4.1 independently of the rank of $\check{R}$. The periodic observability condition (Proposition 2) is the system-theoretic condition that guarantees LMI feasibility with a non-degenerate solution. The constraint $\check{X}\succeq\epsilon I_{Nn}$ (LMI #2) ensures that the Lyapunov matrix is bounded away from singularity, which is a numerical safeguard for solver robustness.
@@ 4.1.5 Stability Guarantee
  5. 5.
    Guaranteed stability: Schur stability is ensured by the Lyapunov inequality structure (Theorem 4.1)
Note: To validate the proposed LMI solution $\check{X}^{-1}\check{Y}$, I conducted numerical experiments with $\check{R}\succ 0$, where the standard DARE solution exists. Let $\check{K}_{\text{Ric}}$ denote the filter gain obtained from the DARE. The predictor-form gain is then $\check{L}_{\text{Ric}}=\check{A}\check{K}_{\text{Ric}}$. Numerical experiments confirm $\|-\check{X}^{-1}\check{Y}-\check{L}_{\text{Ric}}\|_{F}<10^{-6}$.
Remark 4.2 (Relationship to Classical Kalman Filter Optimality)
The proposed LMI formulation (45) minimizes an upper bound on $\mathrm{trace}(\check{P}_{e})$ rather than $\mathrm{trace}(\check{P}_{e})$ itself, since the Lyapunov inequality $\check{P}_{e}\preceq\check{X}^{-1}$ introduces conservatism. However, when the measurement noise covariance is positive definite ($\check{R}\succ 0$), the LMI solution coincides with the classical DARE-based Kalman filter gain, as confirmed both theoretically through the Kalman–LQR duality [13] and numerically (the above Note). When $\check{R}$ is only positive semidefinite, the standard DARE is not directly applicable because the inversion of $\check{C}\check{P}\check{C}^{T}+\check{R}$ in (34) fails. In this case, the LMI approach provides the best achievable trace upper bound under the Lyapunov inequality framework, representing a principled relaxation of the classical optimality criterion. The basic LMI design in this section can therefore be regarded as an optimal Kalman filter in the sense of upper-bound minimization, which reduces to the classical Kalman filter when the standard conditions are met.
@@ 5.0.1 Pole Placement Constraints
To guarantee a minimum convergence rate for the estimation error, I constrain the eigenvalues of the error dynamics matrix $\check{A}-\check{L}\check{C}$ to lie within a specified region. For discrete-time systems, a common choice is the disk region:
@@ 5.0.2 Kalman Filter with $l_{2}$-induced Norm Constraint
The LMI (58) admits a scaling degree of freedom: replacing $\check{C}_{z}$ with $\alpha\check{C}_{z}$ and $\gamma^{2}$ with $\alpha^{2}\gamma^{2}$ for any $\alpha>0$ yields an equivalent $l_{2}$-induced norm condition. This scaling is useful when the (2,2) block $\check{X}-\check{C}_{z}^{T}\check{C}_{z}$ causes infeasibility; choosing $\alpha<1$ relaxes the constraint $\check{X}\succ\check{C}_{z}^{T}\check{C}_{z}$, enabling a common Lyapunov matrix $\check{X}$ to satisfy both the Kalman filter LMI (42) and the $l_{2}$-induced norm constraint simultaneously.
@@ 5.0.3 Combined Multi-objective Design
Remark 5.5 (Extension to Polytopic Uncertainty)
The LMI-based framework has the potential for extension to systems with parametric uncertainty. When the system matrix $A$ belongs to a polytope defined by vertices $\{A^{(1)},A^{(2)},\ldots,A^{(M)}\}$, one can pursue a robust filter design by enforcing LMI constraints at each vertex with a common Lyapunov matrix $\check{X}$ and common gain variable $\check{Y}$. However, this extension requires reformulation of the error dynamics, since the nominal cancellation of the control input $u(k)$ in the estimation error equation no longer holds when the system matrix is uncertain. A detailed treatment is left for future work.
@@ 6 Numerical Example: Multirate Automotive State Estimation
To demonstrate the effectiveness of the proposed multirate Kalman filter design method, I consider an automotive navigation system as a practical application. This example illustrates how the cyclic reformulation framework can be applied to a real-world multirate sensing scenario where GPS and wheel speed sensors operate at different sampling rates.
@@ 6.1 System Configuration
I consider an automotive navigation system with the following specifications:
@@ 6.4 Simulation Results
A simulation of 200 time steps (20 seconds) was conducted with initial state $x(0)=[0,5,0]^{T}$ (position 0 m, velocity 5 m/s), control input $u(k)=0.5\sin(0.05k)$ (sinusoidal acceleration), and process and measurement noise drawn from zero-mean Gaussian distributions with covariances $Q$ and $R$.
Figs. 13 show a representative estimation trajectory for position, velocity, and acceleration, respectively. Position estimation errors are largest just before GPS updates (Fig. 1) and are corrected sharply when GPS measurements become available.
To statistically validate the filter performance, Monte Carlo simulations with 500 independent noise realizations were conducted. Each run uses the same system parameters and initial conditions, with independently generated process and measurement noise sequences. The RMSE for each run is computed over the steady-state interval (time steps 51–200), discarding the first 50 steps as transient. This cutoff is based on the maximum closed-loop eigenvalue magnitude of 0.967, corresponding to a time constant of approximately 30 steps.
Table 2: Monte Carlo Simulation Results (500 runs, steady-state evaluation)
State RMSE 95% Pctl Theory UB Ratio
Position 0.561 m [0.317, 0.919] 0.587 m 1.046
Velocity 0.431 m/s [0.372, 0.490] 0.432 m/s 1.002
Accel. 1.125 m/s2 [0.889, 1.380] 1.130 m/s2 1.005
Table 2 summarizes the Monte Carlo results. The column “Theory UB” lists the theoretical steady-state RMSE for each state, and the column “Ratio” is defined as Theory UB divided by the Monte Carlo mean RMSE. Because the cyclic error covariance upper bound $\check{X}^{-1}\in\mathbb{R}^{Nn\times Nn}$ contains $N$ diagonal blocks of size $n\times n$, one for each time step within the frame period, the theoretical RMSE for state $j$ is computed by averaging the corresponding diagonal entries over the $N$ time steps:
$\text{Theory UB}_{j}=\sqrt{\frac{1}{N}\sum_{k=0}^{N-1}[\check{X}^{-1}]_{(kn+j,\,kn+j)}}$ (72)
This period-averaged value provides a fair comparison with the Monte Carlo mean RMSE, which is also computed over all time steps in the steady-state interval. Since the LMI formulation guarantees $\check{P}_{e}\preceq\check{X}^{-1}$ (see (44)), the theoretical values are guaranteed to be no less than the true steady-state RMSE. The Monte Carlo mean RMSE converges to this true value as the number of runs increases; therefore, the ratio exceeding unity for all three states confirms that the LMI solution provides a valid and consistent upper bound. The ratios are close to unity (1.002–1.046), indicating that the bound is tight with low conservatism.
Regarding estimation accuracy, the position RMSE (0.561 m) is well below the GPS noise standard deviation (1.0 m), confirming that multirate sensor fusion effectively reduces position estimation error. The velocity RMSE (0.431 m/s) exceeds the wheel speed sensor noise standard deviation (0.316 m/s); this is expected because the dominant error source for velocity estimation is not measurement noise but the propagation of acceleration process noise ($Q_{33}=0.5$), which is consistent with the theoretical covariance analysis.